# Control Barrier Functions via Minkowski Operations for Safe Navigation among Polytopic Sets

### v2

@@ VI Conclusion and Future Work
This paper presents a new CBF framework based on the exact SDF between polygons, obtained via convex optimization in $\mathcal{M}_{\mathcal{D}}$-space. We have shown that, in the minimum distance case, the proposed CBF is differentiable a.e. and always has a well-defined subdifferential. For penetration depth, we empirically show that the proposed optimization reformulation is compatible with our CBF framework. We validate our approach in three scenarios that highlight non-conservative maneuvering, collision recovery, and applicability to multi-obstacle environment. Future work includes addressing the limitations of numerically approximating the derivatives of CO bounding constraints by deriving analytical expressions and developing a formal proof and non-smooth analysis for the penetration depth formulation.
@@ ACKNOWLEDGMENT
We thank David Mount at the University of Maryland for his valuable insights on Minkowski operation, and Akshay Thirugnanam at the University of California, Berkeley for helpful discussions on robot parameterization.

### v3

@@ VI Conclusion and Future Work
This paper presents a new CBF framework based on the exact SDF between polygons, obtained via convex optimization in $\mathcal{M}_{\mathcal{D}}$-space. We have shown that, in the minimum distance case, the proposed CBF is differentiable a.e. and always has a well-defined subdifferential. For penetration depth, we empirically show that the proposed optimization reformulation is compatible with our CBF framework. We validate our approach in three scenarios that highlight non-conservative maneuvering, collision recovery, and applicability to multi-obstacle environment. Future work includes deriving analytical expressions of CO bounding constraints for rotational cases and performing a formal nonsmooth analysis for the penetration depth case.
@@ ACKNOWLEDGMENT
We thank David Mount at the University of Maryland for his valuable insights on Minkowski operation, and Akshay Thirugnanam at the University of California, Berkeley for helpful discussions on robot parameterization. This work was partially funded by the NSF under grant NSF GCR 2219101.